Activity 1.c - Identifying Differences between Cash and Accrual Basis

Read each scenario, decide whether the company is using Cash basis or Accrual basis, and then enter your answer to the question.

The Johnny Flowers Law Firm prepays for advertising in the local newspaper. On January 1, the law firm paid $180 for six months of advertising. Johnny Flowers Law Firm recorded $180 in the Prepaid Advertising account.

Accrual basisCash basiscorrect

If Johnny Flowers Law Firm had recorded their expenses using the other method, how much advertising expense would they have recorded for the two months ending February 28? Enter this value as a positive number.

correct

Secure Home provides house-sitting for people while they are away on vacation. Some of its customers pay immediately after the job is finished. Some customers ask that the business send them a bill. As of the end of the year, Secure Home has collected $2,130 from cash-paying customers. Secure Home’s remaining customers owe the business $3,390. Secure Home recorded $2,130 of service revenue for the year.

Accrual basisCash basiscorrect

If Secure Home had recorded their service revenue using the other method, how much service revenue would they have recorded for the year?

correct

Splendid Occasions received $430 for services to be performed for the next 8 months on March 31 and recorded this transaction using the Unearned Revenue account.

Accrual basisCash basiscorrect

If Splendid Occasions had recorded their service revenue using the other method, how much service revenue would they have recorded for the year?

correct

Sweet Catering completed the following selected transactions during May 2016:

  • May 1: Prepaid rent for three months, $1,800
  • May 5: Received and paid electricity bill, $230
  • May 9: Received cash for meals served to customers, $1,040
  • May 14: Paid cash for kitchen equipment, $3,610
  • May 23: Served a banquet on account, $2,640
  • May 31: Made the adjusting entry for rent (from May 1).
  • May 31: Accrued salary expense, $2,430
  • May 31: Recorded depreciation for May on kitchen equipment, $310

If Sweet Catering had recorded transactions using the Cash method, how much net income (loss) would they have recorded for the month of May? If there is a loss, enter it with parentheses or a negative sign.

incorrect

If Sweet Catering had recorded transactions using the Accrual method, how much net income (loss) would they have recorded for the month of May? If there is a loss, enter it with parentheses or a negative sign.
Answer #1

Part 1

If Sweet Catering had recorded transactions using the Cash method, how much net income (loss) would they have recorded for the month of May?

Net income (loss) using cash method = meals served to customer – rent paid – electricity bill – cash paid for kitchen equipment = 1040-1800-230-3610 = $(4600)

Part 2

If Sweet Catering had recorded transactions using the Accrual method, how much net income (loss) would they have recorded for the month of May? If there is a loss, enter it with parentheses or a negative sign.

Net income (loss) using accrual method = meals served to customer+ Served a banquet on account – rent expense – electricity bill – salary expense – depreciation = 1040+2640-(1800/3)-230-2430-310 = $110